# Beyond Winning Strategies: Admissible and Admissible Winning Strategies for Quantitative Reachability Games

### v1

@@ 5 Examples
We now discuss the emergent behavior under admissible and admissible winning strategies in two settings: (i) a gridworld domain, where two agents take actions in turns, and (ii) a manipulator domain, where a robotic arm and a human operate in a shared workspace. For both domains, the task $\varphi$ is specified using $\textsc{ltl}_{f}$ formulas. We first construct a game abstraction $\mathcal{G}$. Next, we construct a deterministic finite automaton (DFA) for the task (Fuggitti 2019) and take the product to construct the product (DFA) game, where the objective for the Sys player is to reach a set of goal (accepting) states. See (Muvvala, Amorese, and Lahijanian 2022; Watanabe et al. 2021) for more details. Code is included in the supplementary material and will be made public upon publication.
Figure 4: Left: Initial setup. Right: Strategy $\sigma^{win}_{adm}$ (red) and $\sigma_{adm}$ (solid white).
Gridworld
@@
Fig. 3 illustrates a gridworld domain with Sys (red), Env (blue), goal (green), and lava (orange) states. The objective for Sys is to reach the goal state while avoiding the Env player. The players should not enter lava states and the Env player can not traverse through the goal state. Action cost is 1 for all Sys player actions. The game is played in turns starting with Sys player where both players take one step in each cardinal direction per turn.
In Fig. 3, dashed line represents dominant strategy for Sys player from current state and solid line represents past actions. Fig. 3(a) and 3(d) illustrate the initial position of both players for two scenarios, 3(b) & 3(e) shows $\sigma_{adm}$, and 3(c) & 3(f) show $\sigma^{win}_{adm}$ strategy. In Fig. 3(a), a winning strategy exists if $\mathcal{B}\geq 12$ but in Fig. 3(d) no winning strategy exists. In Fig. 3(b), for $\mathcal{B}=5$, $\sigma_{adm}$ exists and it relies on Env’s cooperation to reach goal state. Note that $\sigma^{win}_{adm}\equiv\sigma_{adm}$ as there is no winning region. In Fig. 3(c), $\sigma^{win}_{adm}$ exists and commits to go around and reach goal state while $\sigma_{adm}$ strategy does not. For Fig. 3(d) - 3(f), given history, the Sys player is in $V_{win}$. Under strategy $\sigma_{adm}$ in 3(e), the Sys player can go west (in $V_{pen}$) or north (in $V_{win}$). Nonetheless, $\sigma^{win}_{adm}$ in 3(f) commits to going north as it is winning, and there also exists a shorter safer path (pink) to goal state if Env cooperates.
Manipulator Domain
@@
This domain considers a manipulator arm (Sys) operating in presence of a human (Env). In Fig. 4, only Sys can manipulate boxes at location “L#” (bottom) while both players can moves boxes at “HL#” (right) locations. Fig. 4-left shows the initial setup: $b_{0}$ at L0 and $b_{1}$ and HL3. Objective for Sys is: either $b_{0}$ at L1 or $b_{0}$ at HL3 and $b_{1}$ at HL2 and never $b_{0}$ at HL0. It is more expensive for Sys to operate at L# locations than to operate at HL# locations. Note that a winning strategy from the initial state exists.
@@
6 Conclusion
This paper relaxes the requirement of winning strategies in quantitative, reachability games using the notion of admissibility. While we show that admissible strategies are desirable in such settings, they are hard to synthesize due to their history-dependence. We show that such strategies can produce overly optimistic behaviors, and propose admissible winning strategies to mitigate them. We specifically show that admissible winning strategies are appropriate for robotics applications and their synthesis does not require more effort than synthesizing admissible strategies. Future work should explore derivates of admissible winning strategies that are more risk-averse, like regret-minimizing admissible strategies.
Appendix A Appendix
Derivation of Eq. (7)
@@ Appendix A Appendix
A.1 Proof of Lemma 1
@@ A.2 Proof of Lemma 2
We can repeat this for all histories $h$ in $\operatorname{Plays}^{v_{0}}(\sigma)$. Hence, every $\sigma$ that is $\operatorname{\mathbf{SC}}$ is admissible.
@@ A.3 Proof of Theorem 1
[…]
Figure 6: Illustrative example $\mathcal{G}$: for all $h\in\operatorname{Plays}^{v_{0}}$ we define $\operatorname{Val}=0$ if $\operatorname{last}(h)$ is not leaf node else $\operatorname{Val}>0$. The payoffs associated with plays are shown along the edges in black. The values in blue and red are $\operatorname{cVal}(v)$ and $\operatorname{aVal}(v)$.
To prove this it is sufficient to show an example. Consider the game $\mathcal{G}$ in Fig. 6. Let us consider strategies: $\sigma_{1}:(v_{0}\to v_{1}),(v_{3}\to v_{6}),(v_{7}\to v_{8})$, $\sigma_{2}:(v_{0}\to v_{1}),(v_{3}\to v_{6}),(v_{7}\to v_{9})$, $\sigma_{3}:(v_{0}\to v_{2}),(v_{4}\to v_{6}),(v_{7}\to v_{9})$, and $\sigma_{4}:(v_{0}\to v_{2}),(v_{4}\to v_{6}),(v_{7}\to v_{8})$. Note, $\sigma_{1},\sigma_{2}$ and $\sigma_{3},\sigma_{4}$ only differ at state $v_{7}$. Further, $\mathcal{G}$ is a tree-like arena where we define payoff as follows, if $\operatorname{last}(h)$ is a leaf node, then $\operatorname{Val}(\operatorname{last}(h))>0$ else, it is 0. Thus, the payoff function is history-independent, i.e., $\operatorname{Val}(h)=\operatorname{Val}(\operatorname{last}(h))$. Hence, it is sufficient to look at the last state along $h$ to compute optimal adversarial and cooperative values, i.e., $\operatorname{aVal}(h)=\operatorname{aVal}(\operatorname{last}(h))$ and $\operatorname{cVal}(h)=\operatorname{cVal}(\operatorname{last}(h))$.
A.4 Proof of Lemma 3
Proof.
Figure 7: $\operatorname{\mathbf{WCo-Op}}$ proof example
Case I
Case II
A.5 Proof of Theorem 2
Proof.
We begin by showing that memoryless strategies are sufficient for optimal $\operatorname{cVal}$ and $\operatorname{aVal}$. For $\operatorname{\mathbf{Co-Op}}$ strategies, the game can be viewed as a single-player as both players are playing cooperatively. Thus, synthesizing a witnessing strategy $\sigma$ for $\operatorname{\mathbf{Co-Op}}$ reduces to the classical shortest path problem in a weighted graph. This can be solved in polynomial time using Dijkstra’s and Flyod Warshall’s algorithms when the weights are non-negative and arbitrary (Floyd 1962; Mehlhorn and Sanders 2008).
In the adversarial setting, Khachiyan et al. (2008) show that memoryless strategies are sufficient for two-player, non-negative weights scenario. Thus, we have that $\operatorname{cVal}(h,\sigma)=\operatorname{cVal}(\operatorname{last}(h),\sigma)$ for $\operatorname{\mathbf{Co-Op}}$ and $\operatorname{aVal}(h,\sigma)=\operatorname{aVal}(\operatorname{last}(h),\sigma)$ for $\operatorname{\mathbf{WCO}}$.
@@ A.8 Proof of Theorem 4
The result follows from Thm. 7 that states that $\operatorname{\mathbf{SC}}$ or $\operatorname{\mathbf{WCo-Op}}$ are necessary and sufficient conditions for a strategy to be admissible. Using the results of (Brihaye et al. 2017, Thm. 1), we show that the value iteration algorithm on $\mathcal{G}$ will converge after finite iterations to a fixed point. Then, using (Brihaye et al. 2017, Corollary 18 and Proposition 19), we prove that witnessing strategies for $\operatorname{\mathbf{WCO}}$ and $\operatorname{\mathbf{Co-Op}}$ always exists.
@@ A.10 Proof of Lemma 5
While it is sufficient to look only at the last state along a history $h$ to compute $\operatorname{aVal}$, for $\operatorname{\mathbf{SC}}$ condition, $\operatorname{cVal}(v^{\prime})<\operatorname{aVal}(\operatorname{last}(h))$ does not imply that $\operatorname{cVal}(v^{\prime})<\operatorname{aVal}(h_{\leq j})\;\forall j\in\mathbb{N}_{|h|-1}$. This is also evident from Figure 6 where the adversarial values for Sys player states along $h:=v_{0}v_{1}v_{3}v_{6}v_{7}$ and $v^{\prime}:=v_{9}$ is $\operatorname{aVal}(v_{0})=\infty;\operatorname{aVal}(v_{3})=4;\operatorname{aVal}(v_{7})=9$. While $\operatorname{cVal}(v_{9})<\operatorname{aVal}(v_{7})$, $\operatorname{cVal}(v_{9})>\operatorname{aVal}(v_{3})$.
@@ A.11 Proof of Theorem 5
Let us define $\operatorname{aVal}(\mathcal{G},v)$ and $\operatorname{cVal}(\mathcal{G},v)$ to be the optional adversarial and cooperative value for state $v$ in $\mathcal{G}$. We extend the definition to $\operatorname{acVal}(\mathcal{G},v)$ accordingly. We begin our proof by first noting that $\mathcal{G}^{\prime}$ is a finite tree arena where the leaf nodes are partitioned into goal states for the Sys player and sink states. By construction, every state that is not a leaf node in $\mathcal{G}^{\prime}$ has at least one outgoing edge. Only plays that reach a goal state in $\mathcal{G}^{\prime}$ have a finite payoff value. Plays that fail to reach a goal state have a payoff of +$\infty$+ ∞. Thus, using Khachiyan et al. (2008)’s algorithm, we can compute optimal $\operatorname{aVal}(\mathcal{G}^{\prime},v)$ and $\operatorname{cVal}(\mathcal{G}^{\prime},v)$ and the corresponding witnessing strategies. Thus, game $\mathcal{G}^{\prime}$ is well-formed, which implies $\operatorname{\mathbf{SC}}$ and $\operatorname{\mathbf{WCo-Op}}$ always exists (from Thm. 4). Hence, there always exists an admissible strategy in $\mathcal{G}^{\prime}$.
@@ A.12 Proof of Lemma 6
Given $\mathcal{G}$, the initial state of the game can belong to three regions. Thus, $v_{0}$ is either part of the losing, pending, or winning region. From Thm. 7, a strategy $\sigma$ is admissible if, and only if, it is $\operatorname{\mathbf{SC}}$ or $\operatorname{\mathbf{WCo-Op}}$.
@@ A.13 Proof of Theorem 10
Now we use the same techniques from Thm. 7 to prove that $\operatorname{\mathbf{mSC}}$ and $\operatorname{\mathbf{WCo-Op}}$ are necessary and sufficient conditions for a strategy to be admissible winning.
Proof of Thm. 10.
@@ A.14 Proof of Lemma 7
From Thm. 10 we have that $\operatorname{\mathbf{WCo-Op}}$ is a necessary and sufficient condition for a strategy to be admissible winning. From Thm. 2, we have that $\operatorname{\mathbf{WCo-Op}}$ always exists. Thus, we conclude that admissible winning strategies always exist.
Further, we can show that $\operatorname{\mathbf{mSC}}$ strategies always exist. To prove this, we observe that $\operatorname{\mathbf{mSC}}$ is a subset of $\operatorname{\mathbf{SC}}$. Specifically, $\forall h\in\operatorname{Plays}^{h}(\sigma)$ such that $\operatorname{aVal}(h)<\infty$, $\operatorname{\mathbf{mSC}}$ are exactly the strategies that are $\operatorname{\mathbf{SC}}$ and value-preserving. From Lemma 8, we have that $\operatorname{\mathbf{mSC}}$ are winning strategies for all $h$ such that $\operatorname{aVal}(h)<\infty$. Thus, witnessing strategies for worst-case strategies from the winning region is a sufficient condition for $\operatorname{\mathbf{mSC}}$ to always exist. From (Brihaye et al. 2017) and A.9, we can compute witnessing strategies for all worst-case values for every history $h$ in $\mathcal{G}$. Thus, $\operatorname{\mathbf{mSC}}$ strategies always exist. ∎
A.15 Proof of Theorem 7

### v2

@@ 5 Illustrative Examples
We now discuss the emergent behavior under admissible and admissible winning strategies in two settings: (i) a gridworld domain, where two agents take actions in turns, and (ii) a manipulator domain, where a robotic arm and a human operate in a shared workspace. For both domains, the task $\varphi$ is specified using $\textsc{ltl}_{f}$ formulas. We first construct a game abstraction $\mathcal{G}$. Next, we construct a deterministic finite automaton (DFA) for the task [Fuggitti, 2019] and take the product to construct the product (DFA) game, where the objective for the Sys player is to reach a set of goal (accepting) states. Code is included in the supplementary material and will be made public upon publication. Additionally, more manipulator domain experiments and empirical validations of PTIME complexity of Alg. 1 for fixed $\mathcal{B}$ are provided in the Appendix. Videos for all experiments are provided in the supplementary material.
(a) $\sigma_{adm},5$
(b) $\sigma^{win}_{adm},12$
(c) $\sigma_{adm},8$
(d) $\sigma^{win}_{adm},8$
Figure 3: Gridworld Example. Sub-captions show strategy and $\mathcal{B}$.
Gridworld.
@@
Fig. 3 illustrates a gridworld domain with Sys (red), Env (blue), goal (green), and lava (orange) states. The objective for Sys is to reach the goal state while avoiding the Env player. The players should not enter lava and Env player cannot traverse through goal state. Action cost is 1 for all Sys player actions. The game is played in turns starting with Sys where both players take a step in each cardinal direction. Figs. 3(a) and 3(c) illustrate the initial position of both players for two different scenarios. Figs. 3(a)-3(b) and 3(c)-3(d) illustrate $\sigma_{adm}$ and $\sigma^{win}_{adm}$ for different budgets. In Figs. 3(a)-3(b), a winning strategy exists if $\mathcal{B}\geq 12$. Hence, for $\mathcal{B}=5$ (Fig. 3(a)), only $\sigma_{adm}$ exists, which relies on Env’s cooperation to reach goal state. For $B=12$ (Fig. 3(b)), $\sigma^{win}_{adm}$ exists and commits to go around and reach the goal state. Note that $\sigma_{adm}$ (not shown) also exists but does not commit to go around.
In Fig. 3(c), we start in the winning region. Despite $v_{0}\in V_{win}$, $\sigma_{adm}$ chooses to go west (in $V_{pen}$) or north (in $V_{win}$). In contrast, $\sigma^{win}_{adm}$ in Fig. 3(d) stays in the winning region (goes north) as it is the sole winning strategy that is also admissible, with a potentially shorter path (dashed) if Env cooperates.
Comparison against Best-Effort (BE): Here we illustrate differences between BE and admissible strategies. Besides qualitative vs quantitative, in the winning region, a major difference is that every $\sigma_{win}$ is also BE, but not every $\sigma_{win}$ is winning admissible. In Fig. 3(d), if Env player stays in the left corner (cooperative), $\sigma^{win}_{adm}$ commits to dashed-line strategy as the solid-line strategy is not $\operatorname{\mathbf{Co-Op}}$, whereas BE does not distinguish between the two strategies as both are winning.
Figure 4: Left: Initial setup. Right: $\sigma_{adm}$ for $\mathcal{B}=13$. Task: $\varphi=\bigl{(}p_{04}\wedge p_{15}\wedge\raisebox{-1.04625pt}{\LARGE$\circ$}(p_{26}\vee p_{36})\bigr{)}$ where $p_{ij}$ is block $i$ at loc $j$. Blocks 0 to 2 are boxes and block 3 is a cylinder.
Manipulator Domain.
This domain considers a robotic arm (Sys) operating in presence of a human (Env), as described in [Muvvala et al., 2022]. The task for the robot is to build an arch with two boxes as support and one block (cylinder or box) on top as shown in Fig. 4. The human can choose to either move a block back to the initial position or not intervene. The robot’s actions are transit, grasp, transfer, and release with unit cost for each action, except for transferring the cylinder, which costs two units.
A winning strategy from the initial state does not exist as the human can always undo robot actions. However, an admissible strategy $\sigma_{adm}$ that solves Problem 1 exists. Here, $\sigma_{adm}$ continually attempts to build the arch with either a box or the cylinder on top. Interestingly, a subset of $\sigma_{adm}$, $\operatorname{\mathbf{WCo-Op}}$ strategies, commit to building an arch with the box on top as it ensures the least payoff among all $\operatorname{\mathbf{WCO}}$ strategies. Note that synthesizing $\operatorname{\mathbf{WCo-Op}}$ is easy as it can be constructed from $\operatorname{\mathbf{WCO}}$ and $\operatorname{\mathbf{Co-Op}}$ (Line 2 of Alg. 1). Thus, $\sigma_{adm}$ allows the robot to operate beyond the winning region while more desirable (less costly) behaviors can also be extracted via $\operatorname{\mathbf{WCo-Op}}$ strategies.
Comparison against Best-Effort (BE): BE strategies build arch with either a box or cylinder on top, whereas $\operatorname{\mathbf{WCo-Op}}$ strategies use a box on top, which is more desirable.
6 Conclusion
This paper relaxes the requirement of winning strategies in quantitative, reachability games using the notion of admissibility. While we show that admissible strategies are desirable in such settings, they are hard to synthesize due to their history-dependence. We show that such strategies can produce overly optimistic behaviors, and propose admissible winning strategies to mitigate them. We specifically show that admissible winning strategies are appropriate for robotics applications and their synthesis does not require more effort than synthesizing admissible strategies. Future work should explore classes of admissible winning strategies that are more risk-averse, e.g., regret-minimizing admissible strategies.
Appendix A Appendix - Additional Experiments
Figure 5: Left: Initial setup. Right: Strategy $\sigma^{win}_{adm}$ (red) and $\sigma_{adm}$ (solid white). $\varphi=(\neg p_{09})\mathop{\mathcal{U}}(p_{01}\vee(p_{06}\wedge p_{17}))$, where $p_{ij}$ is box $i$ in loc $j$ where loc 9 is HL0; loc 6 & 7 are HL3 & HL2, respectively.
Manipulator Example 2.
This domain again considers a robotic arm (Sys) operating in presence of a human (Env). Similar to the manipulator example considered in the Results section. In this example, we illustrate the difference in the emergent behavior when an admissible winning strategy as well as an admissible strategy exists.
@@ Appendix A Appendix - Additional Experiments
In Fig. 5, only Sys can manipulate boxes at location “L#” (bottom) while both players can moves boxes at “HL#” (right) locations. Fig. 5-left shows the initial setup: $b_{0}$ at L0 and $b_{1}$ and HL3. Objective for Sys is: either $b_{0}$ at L1 or $b_{0}$ at HL3 and $b_{1}$ at HL2 and never $b_{0}$ at HL0. It is more expensive for Sys to operate at L# locations than to operate at HL# locations. Note that a winning strategy from the initial state exists.
@@
Computation Results.
In Fig. 6, we report computation times that empirically shows PTIME complexity for Alg. 1 for fixed $\mathcal{B}$ for the Manipulator case-study. We consider 4 different specifications: $\varphi_{1}=F(p_{06})$; $\varphi_{2}=F(p_{06}\wedge F(p_{07}))$; $\varphi_{3}=F(p_{06}\wedge F(p_{07}\wedge F(p_{08})))$; $\varphi_{4}=F(p_{06}\wedge F(p_{07}\wedge F(p_{08}\wedge F(p_{09}))))$.
Appendix B Appendix - Proofs and Derivation
Derivation of Eq. (7).
@@ Appendix B Appendix - Proofs and Derivation
B.1 Necessary and Sufficient Conditions for weakly dominating strategy
A strategy $\sigma$ is not admissible if there exists another strategy $\sigma^{\prime}$ that weakly dominates it. Formally, $\sigma^{\prime}\succ\sigma$ if, and only if, $\forall h\in\operatorname{Plays}^{v_{0}}(\sigma)$, the following holds:
Lemma 7.
Figure 6: PTIME complexity for Alg. 1 for fixed Computation Time vs Budget for fixed $\mathcal{G}$ and varying $\varphi$. See Section A for details.
B.2 Proof of Lemma 1
@@ B.3 Proof of Lemma 2
We can repeat this for all histories $h$ in $\operatorname{Plays}^{v_{0}}(\sigma)$. Hence, every $\sigma$ that is $\operatorname{\mathbf{SC}}$ is admissible. ∎
B.4 Proof of Theorem 1
Proof.
To prove this it is sufficient to show an example. Consider the game $\mathcal{G}$ in Fig. 8. Let us consider strategies: $\sigma_{1}:(v_{0}\to v_{1}),(v_{3}\to v_{6}),(v_{7}\to v_{8})$, $\sigma_{2}:(v_{0}\to v_{1}),(v_{3}\to v_{6}),(v_{7}\to v_{9})$, $\sigma_{3}:(v_{0}\to v_{2}),(v_{4}\to v_{6}),(v_{7}\to v_{9})$, and $\sigma_{4}:(v_{0}\to v_{2}),(v_{4}\to v_{6}),(v_{7}\to v_{8})$. Note, $\sigma_{1},\sigma_{2}$ and $\sigma_{3},\sigma_{4}$ only differ at state $v_{7}$ and $v_{0}$.
[…]
Figure 8: Illustrative example $\mathcal{G}$: for all $h\in\operatorname{Plays}^{v_{0}}$ we define $\operatorname{Val}<\infty$ if a play reaches a goal (double circle around the state) state. The cost of the Env action is zero, while the cost of the Sys actions are shown along the edges. The values in blue and red are $\operatorname{cVal}(v)$ and $\operatorname{aVal}(v)$.
B.5 Proof of Lemma 3
Proof.
Figure 9: $\operatorname{\mathbf{WCo-Op}}$ proof example
Case I
Case II
@@ B.6 Proof of Theorem 2
We begin by showing that memoryless strategies are sufficient for optimal $\operatorname{cVal}$ and $\operatorname{aVal}$. For $\operatorname{\mathbf{Co-Op}}$ strategies, the game can be viewed as a single-player as both players are playing cooperatively. Thus, synthesizing a witnessing strategy $\sigma$ for $\operatorname{\mathbf{Co-Op}}$ reduces to the classical shortest path problem in a weighted graph. This can be solved in polynomial time using Dijkstra’s and Flyod Warshall’s algorithms when the weights are non-negative and arbitrary Floyd [1962]; Mehlhorn and Sanders [2008].
In the adversarial setting, Khachiyan et al. [2008] show that memoryless strategies are sufficient for two-player, non-negative weights scenario. Thus, we have that $\operatorname{cVal}(h,\sigma)=\operatorname{cVal}(\operatorname{last}(h),\sigma)$ for $\operatorname{\mathbf{Co-Op}}$ and $\operatorname{aVal}(h,\sigma)=\operatorname{aVal}(\operatorname{last}(h),\sigma)$ for $\operatorname{\mathbf{WCO}}$.
@@ B.9 Proof of Theorem 4
The result follows from Thm. 3 that states that $\operatorname{\mathbf{SC}}$ or $\operatorname{\mathbf{WCo-Op}}$ are necessary and sufficient conditions for a strategy to be admissible. Using the results of [Brihaye et al., 2017, Thm. 1], we show that the value iteration algorithm on $\mathcal{G}$ will converge after finite iterations to a fixed point. Then, using [Brihaye et al., 2017, Corollary 18 and Proposition 19], we prove that witnessing strategies for $\operatorname{\mathbf{WCO}}$ and $\operatorname{\mathbf{Co-Op}}$ always exists.
@@ B.11 Proof of Lemma 4
While it is sufficient to look only at the last state along a history $h$ to compute $\operatorname{aVal}$, for $\operatorname{\mathbf{SC}}$ condition, $\operatorname{cVal}(v^{\prime})<\operatorname{aVal}(\operatorname{last}(h))$ does not imply that $\operatorname{cVal}(v^{\prime})<\operatorname{aVal}(h_{\leq j})\;\forall j\in\mathbb{N}_{|h|-1}$. This is also evident from Figure 8 where the adversarial values for Sys player states along $h:=v_{0}v_{1}v_{3}v_{6}v_{7}$ and $v^{\prime}:=v_{9}$ is $\operatorname{aVal}(v_{0})=\infty;\operatorname{aVal}(v_{3})=6;\operatorname{aVal}(v_{7})=10$. While $\operatorname{cVal}(v_{9})<\operatorname{aVal}(v_{7})$, $\operatorname{cVal}(v_{9})>\operatorname{aVal}(v_{3})$.
@@ B.12 Proof of Theorem 5
Let us define $\operatorname{aVal}(\mathcal{G},v)$ and $\operatorname{cVal}(\mathcal{G},v)$ to be the positional adversarial and cooperative value for state $v$ in $\mathcal{G}$. We extend the definition to $\operatorname{acVal}(\mathcal{G},v)$ accordingly. We begin our proof by first noting that $\mathcal{G}^{\prime}$ is a finite tree arena where the leaf nodes are partitioned into goal states for the Sys player and sink states. By construction, every state that is not a leaf node in $\mathcal{G}^{\prime}$ has at least one outgoing edge. Only plays that reach a goal state in $\mathcal{G}^{\prime}$ have a finite payoff value. Plays that fail to reach a goal state have a payoff of +$\infty$+ ∞. Thus, using Khachiyan et al. [2008]’s algorithm, we can compute optimal $\operatorname{aVal}(\mathcal{G}^{\prime},v)$ and $\operatorname{cVal}(\mathcal{G}^{\prime},v)$ and the corresponding witnessing strategies. Thus, game $\mathcal{G}^{\prime}$ is well-formed, which implies $\operatorname{\mathbf{SC}}$ and $\operatorname{\mathbf{WCo-Op}}$ always exists (from Thm. 4). Hence, there always exists an admissible strategy in $\mathcal{G}^{\prime}$.
@@ B.13 Proof of Lemma 5
Given $\mathcal{G}$, the initial state of the game can belong to three regions. Thus, $v_{0}$ is either part of the losing, pending, or winning region. From Thm. 3, a strategy $\sigma$ is admissible if, and only if, it is $\operatorname{\mathbf{SC}}$ or $\operatorname{\mathbf{WCo-Op}}$.
@@ B.14 Proof of Theorem 6
Now we use the same techniques from Thm. 3 to prove that $\operatorname{\mathbf{mSC}}$ and $\operatorname{\mathbf{WCo-Op}}$ are necessary and sufficient conditions for a strategy to be admissible winning.
Proof of Thm. 6.
@@ B.15 Proof of Lemma 6
From Thm. 6 we have that $\operatorname{\mathbf{WCo-Op}}$ is a necessary and sufficient condition for a strategy to be admissible winning. From Thm. 2, we have that $\operatorname{\mathbf{WCo-Op}}$ always exists. Thus, we conclude that admissible winning strategies always exist.
Further, we can show that $\operatorname{\mathbf{mSC}}$ strategies always exist. To prove this, we observe that $\operatorname{\mathbf{mSC}}$ is a subset of $\operatorname{\mathbf{SC}}$. Specifically, $\forall h\in\operatorname{Plays}^{h}(\sigma)$ such that $\operatorname{aVal}(h)<\infty$, $\operatorname{\mathbf{mSC}}$ are exactly the strategies that are $\operatorname{\mathbf{SC}}$ and value-preserving. From Lemma 8, we have that $\operatorname{\mathbf{mSC}}$ are winning strategies for all $h$ such that $\operatorname{aVal}(h)<\infty$. Thus, witnessing strategies for worst-case strategies from the winning region is a sufficient condition for $\operatorname{\mathbf{mSC}}$ to always exist. From Brihaye et al. [2017] and B.10, we can compute witnessing strategies for all worst-case values for every history $h$ in $\mathcal{G}$. Thus, $\operatorname{\mathbf{mSC}}$ strategies always exist. ∎
B.16 Proof of Theorem 7